OpenStudy (anonymous):

Will fan and Medal who ever can help? A pharmacist wants to make 20 ml of a 59% hydrogen peroxide solution. How many mls of 35% hydrogen peroxide must he mix with 75% hydrogen peroxide to make the desired solution?

OpenStudy (mathmate):

Ratio of vol of 35% : vol of 75% =(75-59) : (59-35) =16: 24 Vol of 35% required = 16/(16+24)* 20 ml =8 ml Vol of 75% required = 24/(16+24)* 20 ml = 12 ml. You can also solve this by setting up a system of 2 equations and solving it.

OpenStudy (anonymous):

so the answer is 8 of 35% and 75% of 12

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Yes, that is correct. But do you know how it is figured out?

OpenStudy (anonymous):

yea you showed me

OpenStudy (mathmate):

Good! Happy to hear that! :)